# 37N35, 65F55, 93A15, 93B15, 93C57

### v1

@@ 4.2 Error control
The result [21, Prop. 3.2] provides error control on the data matrices that appear in $\mathsf{foQuadBT}$ in terms of their intrusive counterparts. By recognizing that the proof of [21, Prop. 3.2] makes no assumption on the underlying system structure, this result can be extended directly to our setting to derive error bounds on the matrices in Theorem 1.
@@
The result follows by carefully following the proof of [21, Prop. 3.2] in the context of Theorem 1. ∎
@@ 5.1 Enforcing real-valued models from complex-valued data
where $\boldsymbol{k}$ and $\boldsymbol{k}+\boldsymbol{1}$ refer to the indices $(k-1)p+1:kp$ and $kp+1:(k+1)p$ and similarly for $\boldsymbol{j}$ and $\boldsymbol{j}+\boldsymbol{1}$. From the above, we observe that $\overline{\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}}=\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}+\boldsymbol{1}}$ and $\overline{\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}}}=\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}}$. Define the unitary matrix $\boldsymbol{J}_{\ell}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\frac{1}{\sqrt{2}}\begin{bmatrix}\boldsymbol{I}_{\ell}&-\mathfrak{i}\boldsymbol{I}_{\ell}\\ \boldsymbol{I}_{\ell}&\hphantom{-}\mathfrak{i}\boldsymbol{I}_{\ell}\end{bmatrix}\in\mathbb{C}^{2\ell\times 2\ell}$ for some positive integer $\ell\in\mathbb{N}$. One can readily verify that
@@
which is real valued. Thus, the matrix $\mathbb{M}$ in eq. 22a is unitarily equivalent to a real-valued matrix $\widetilde{\mathbb{M}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{p}^{\mathsf{H}})\mathbb{M}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{m})\in\mathbb{R}^{pK\times mJ}$, where $\otimes$ denotes the matrix Kronecker product [20, Chap. 12.3]. Similarly, the matrices $\mathbb{K}$, $\mathbb{B}_{\operatorname{u}}$ and $\mathbb{C}_{\operatorname{p}}$ in eqs. 22b and 22c are unitarily equivalent to real-valued matrices as well. Namely, one can verify that
@@
are all real valued by construction. Therefore, we may use these unitary transformations to obtain real-valued data matrices in Algorithm 2 and, consequently, a real-valued realizations of the reduced-order models. Lastly, we mention that this realification procedure can also be applied to the second-order Hermite Loewner matrices derived in Theorem 2. If the underlying system is real valued and the data satisfy the hypotheses of Theorem 3, then this allows one to obtain a real-valued realization of an asymptotically stable system. In this setting, it will be the case that $\zeta_{i+1}=s_{i+1}=-s_{i}=\vartheta_{i}$ and $\zeta_{i}=s_{i}=-s_{i+1}=\vartheta_{i+1}$ due to eq. 30, and so evaluations of the derivative of $\boldsymbol{G}$ will appear in the sub- and super-diagonals of $\mathbb{M}$ and $\mathbb{K}$.
@@ 5.2 Optimal choice of damping parameters
Thus far, we have assumed that the coefficient functions $f(s)$ and $g(s)$ in the generalized proportional damping model eq. 5 are given and can be freely evaluated. However, in a truly data-driven regime, these would be unknown. We discuss how these coefficients can be inferred from data in the simplified cases of Rayleigh damping eq. 6 and structural damping eq. 7; this corresponds to inferring the parameters $\alpha,\beta\in\mathbb{R}$ or $\eta\in\mathbb{R}$, respectively. In either case, we assume the data $\boldsymbol{g}_{k}=\boldsymbol{G}(s_{k})$ to be prescribed at a fixed number of points $s_{k}\in\mathbb{C}$, for $k=1,\ldots,N$; we do not assume this data to have any particular structure other than that it is generated by the true underlying system’s transfer function $\boldsymbol{G}$. The reduced-order matrices are computed according to Algorithm 2, and the goal is then to optimize the damping parameters in the $\mathsf{soQuadpvBT}$ reduced model to best fit the given data. A similar idea has been employed in the previously mentioned second-order Loewner framework [32], as well as in the data-driven modeling of time-delay systems [38]. In contrast to these works, we explicitly derive gradients of the relevant cost function with respect to the damping parameters.
To streamline the subsequent presentation, we introduce the following notation
$\boldsymbol{\varphi}_{\operatorname{R}}$ $\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ with $\boldsymbol{\varphi}_{\operatorname{R}}(s,\alpha,\beta)$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(s^{2}+s\alpha){\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+s\beta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},$
$\boldsymbol{\varphi}_{\operatorname{S}}$ $\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ with $\boldsymbol{\varphi}_{\operatorname{S}}(s,\eta)$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}s^{2}{\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+\mathfrak{i}\eta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},$
$\mathcal{C}$ $\colon\mathbb{C}\to\mathbb{C}^{p\times n}$ with $\mathcal{C}(s)$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}+s{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}.$
Even though $\mathsf{soQuadpvBT}$ constructs reduced models with ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}=\boldsymbol{I}_{r}$, this is not enforced in the subsequent results to make the presentation as general as possible. Thus, the minimization procedure we describe can be applied to any data-driven reduced model eq. 3 with unknown damping parameters; e.g., it can be applied to those computed by the second-order Loewner framework [32]. For the Rayleigh damped case, the optimization problem is formulated as
@@
$\begin{split}(\alpha_{\star},\beta_{\star})&=\operatorname*{arg\,min}\limits_{(\alpha,\beta)\in\mathbb{R}^{2}}\sum\limits_{k=1}^{N}\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta),\quad\text{with}\\ \mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)&\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left\lVert\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\right\rVert_{\operatorname{F}}^{2},\end{split}$ (31)
and for the structurally damped case, we have
@@
$\eta_{\star}=\operatorname*{arg\,min}\limits_{\eta\in\mathbb{R}}\sum\limits_{k=1}^{N}\mathcal{J}_{k,\operatorname{S}}(\eta),\quad\text{with}\quad\mathcal{J}_{k,\operatorname{S}}(\eta)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left\lVert\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\right\rVert_{\operatorname{F}}^{2}.$ (32)
The gradients of the objective functions eqs. 31 and 32 are presented below.
@@
Consider a reduced-order model of the form eq. 3. For the minimization problem eq. 31, the gradients with respect to the optimization parameters are
[…]
@@
[…]
@@
×$\quad{}\times{}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{(}\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{)}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}\bigg{)}\bigg{)}.$
@@
It suffices to show eq. 33b since the proofs of eqs. 33a and 34 follow analogously. Define $\mathcal{E}_{k}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\in\mathbb{C}^{p\times m}$ so that $\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)=\lVert\mathcal{E}_{k}\rVert_{\operatorname{F}}^{2}=\operatorname{tr}(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{E}_{k})$. Now, we consider an arbitrary perturbation $\Delta_{\beta}\in\mathbb{R}$, and expand the helper function $\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta+\Delta_{\beta})$ as
@@
$=\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}-s_{k}\Delta_{\beta}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2}).$
@@
$=\operatorname{tr}(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{E}_{k})+2\operatorname{Re}\Big{(}s_{k}\operatorname{tr}\Big{(}\mathcal{E}_{k}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big{)}\Big{)}\Delta_{\beta}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2})$
$=\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)+2\operatorname{Re}\Big{(}s_{k}\operatorname{tr}\Big{(}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\mathcal{E}_{k}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}\Big{)}\Big{)}\Delta_{\beta}+\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2}),$
@@
In a setting where the damping coefficients are unknown, the gradients in Theorem 5 can be used in any off-the-shelf optimization algorithm to minimize the cost function eq. 31 over possible values of $\alpha$ and $\beta$ simultaneously if the system is Rayleigh damped, or in the case of structural damping, to minimize eq. 32 over $\eta$. This minimization routine adds one additional step at the end of Algorithm 2. We emphasize that the reduced-order model matrices do not need to be recomputed after the optimization step; the computed parameters are only used to form the reduced-order damping term from the already computed mass and stiffness matrices.

### v2

@@ 4.2 Error control
The result [19, Prop. 3.2] provides error control on the data matrices that appear in $\mathsf{foQuadBT}$ in terms of their intrusive counterparts. By recognizing that the proof of [19, Prop. 3.2] makes no assumption on the underlying system structure, this result can be extended directly to our setting to derive error bounds on the matrices in Theorem 1.
@@
Note that the proof of [19, Prop. 3.2] makes no assumptions on the underlying system structure. Thus, the inequalities in eq. 28 follow from carefully executing the proof of [19, Prop. 3.2] in the context of Theorem 1. ∎
@@ 5.1 Enforcing real-valued models from complex-valued data
where $\boldsymbol{k}$ and $\boldsymbol{k}+\boldsymbol{1}$ refer to the indices (k1)p+1:kp(k-1)p+1:kp and $kp+1:(k+1)p$ and similarly for $\boldsymbol{j}$ and $\boldsymbol{j}+\boldsymbol{1}$. From the above, we observe that $\overline{\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}}=\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}+\boldsymbol{1}}$ and $\overline{\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}+\boldsymbol{1}}}=\mathbb{M}_{\boldsymbol{k}+\boldsymbol{1},\boldsymbol{j}}$. Define the unitary matrix $\boldsymbol{J}_{\ell}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\frac{1}{\sqrt{2}}\begin{bmatrix}\boldsymbol{I}_{\ell}&-\mathfrak{i}\boldsymbol{I}_{\ell}\\ \boldsymbol{I}_{\ell}&\hphantom{-}\mathfrak{i}\boldsymbol{I}_{\ell}\end{bmatrix}\in\mathbb{C}^{2\ell\times 2\ell}$ for some positive integer $\ell\in\mathbb{N}$. One can readily verify that
@@
which is real valued. Thus, the matrix $\mathbb{M}$ in eq. 22a is unitarily equivalent to a real-valued matrix $\widetilde{\mathbb{M}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{p}^{\mathsf{H}})\mathbb{M}(\boldsymbol{I}_{N/2}\otimes\boldsymbol{J}_{m})\in\mathbb{R}^{pK\times mJ}$, where $\otimes$ denotes the matrix Kronecker product [18, Chap. 12.3]. Similarly, the matrices $\mathbb{K}$, $\mathbb{B}_{\operatorname{u}}$ and $\mathbb{C}_{\operatorname{p}}$ in eqs. 22b and 22c are unitarily equivalent to real-valued matrices as well. Namely, one can verify that
@@
are all real valued by construction. Therefore, we may use these unitary transformations to obtain real-valued data matrices in Algorithm 2 and, consequently, a real-valued realizations of the reduced-order models. Lastly, we mention that this realification procedure can also be applied to the second-order Hermite Loewner matrices derived in Theorem 2. If the underlying system is real-valued and the data satisfy the hypotheses of Theorem 3, then this allows one to obtain a real-valued realization of an asymptotically stable system. In this setting, it will be the case that $\zeta_{i+1}=s_{i+1}=-s_{i}=\vartheta_{i}$ and $\zeta_{i}=s_{i}=-s_{i+1}=\vartheta_{i+1}$ due to eq. 30, and so evaluations of the derivative of $\boldsymbol{G}$ will appear in the sub- and super-diagonals of $\mathbb{M}$ and $\mathbb{K}$.
@@ 5.2 Optimal choice of damping parameters
Thus far, we have assumed that the coefficient functions $f$ and $g$ in the generalized proportional damping model eq. 5 are given and can be freely evaluated. However, in a truly data-driven regime, these would be unknown. We discuss how these coefficients can be inferred from data in the simplified cases of Rayleigh damping eq. 6 and structural damping eq. 7; this corresponds to inferring the parameters $\alpha,\beta\geq 0$ or $\eta\geq 0$, respectively. In either case, we assume the data $\boldsymbol{g}_{k}=\boldsymbol{G}(s_{k})$ to be prescribed at a fixed number of points $s_{k}\in\mathbb{C}$, for $k=1,\ldots,N$; we do not assume this data to have any particular structure other than that it is generated by the true underlying system’s transfer function $\boldsymbol{G}$. A similar idea has been employed in the previously mentioned second-order Loewner framework [29], as well as in the data-driven modeling of time-delay systems [35].
We estimate the unknown parameters by minimizing a least squares fit of the given data over some realistic range(s) of parameter values, which are typically available in practice. For the Rayleigh-damped case, the optimization problem we aim to solve is
@@
$\begin{split}(\alpha_{\star},\beta_{\star})&=\operatorname*{arg\,min}\limits_{(\alpha,\beta)\in\mathcal{I}_{\alpha}\times\mathcal{I}_{\beta}}\sum\limits_{k=1}^{N}\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta),\quad\text{with}\\ \mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)&\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left\lVert\boldsymbol{g}_{k}-\left({\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}+s_{k}{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\right)\left((s_{k}^{2}+s_{k}\alpha){\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+s_{k}\beta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\right)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\right\rVert_{\operatorname{F}}^{2}\end{split}$ (31)
where $\mathcal{I}_{\alpha}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}[\alpha_{-},\alpha_{+}]$ and $\mathcal{I}_{\beta}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}[\beta_{-},\beta_{+}]$. For the structurally damped case, we have
@@
$\begin{split}\eta_{\star}&=\operatorname*{arg\,min}\limits_{\eta\in\mathcal{I}_{\eta}}\sum\limits_{k=1}^{N}\mathcal{J}_{k,\operatorname{S}}(\eta),\quad\text{with}\\ \mathcal{J}_{k,\operatorname{S}}(\eta)&\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left\lVert\boldsymbol{g}_{k}-\left({\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}+s_{k}{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\right)\left(s_{k}^{2}{\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+\mathfrak{i}\eta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\right)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\right\rVert_{\operatorname{F}}^{2}\end{split}$ (32)
where $\mathcal{I}_{\eta}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}[\eta_{-},\eta_{+}]$. Any suitable off-the-shelf optimization algorithm can be used to minimize the cost function eq. 31 over the possible values of $\alpha\in\mathcal{I}_{\alpha}$ and $\beta\in\mathcal{I}_{\beta}$ simultaneously if the system is Rayleigh damped, or, in the case of structural damping, to minimize eq. 32 over $\eta\in\mathcal{I}_{\eta}$. Even though $\mathsf{soQuadpvBT}$ constructs reduced models with ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}=\boldsymbol{I}_{r}$, this is not enforced here to make the presentation as general as possible. Thus, the minimization procedure we describe can be applied to any data-driven reduced model eq. 3 with unknown damping parameters; e.g., it can be applied to those computed by the second-order Loewner framework [29].
We emphasize that the reduced-order matrices may depend explicitly on the parameters $\alpha$, $\beta$, or $\eta$ as in the case of $\mathsf{soQuadpvBT}$ (Algorithm 2). However, we omit this dependence in our notation to simplify the presentation. In this case, the reduced-order matrices need to be recomputed for every evaluation of the cost function in the minimization procedure. Once the optimal coefficients $\alpha_{\star}\in\mathcal{I}_{\alpha}$ and $\beta_{\star}\in\mathcal{I}_{\beta}$ or $\eta_{\star}\in\mathcal{I}_{\eta}$ are computed by minimizing eq. 31 or eq. 32, respectively, a final reduced model is computed by Algorithm 2 using these values.
If the damping coefficients are treated as fixed in the construction of the reduced model matrices, or if one has access to an already computed reduced model eq. 3, then the damping coefficients are only used to form the reduced-order damping term from the mass and stiffness matrices. In this case, the objective functions eqs. 31 and 32 can be minimized without reassembling the reduced model at every step. To this end, we derive gradients of the objective functions in eqs. 31 and 32 when the reduced-order model matrices do not depend on the damping coefficients. For our purposes, we employ these gradients in an optional post-processing step at the end of Algorithm 2 to form the reduced-order damping matrix.
@@
Consider a reduced-order model of the form eq. 3. Let
$\boldsymbol{\varphi}_{\operatorname{R}}$ $\colon\mathbb{C}\times[0,\infty)\times[0,\infty)\to\mathbb{C}^{n\times n}$ with $\boldsymbol{\varphi}_{\operatorname{R}}(s,\alpha,\beta)$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}(s^{2}+s\alpha){\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+s\beta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}$,
$\boldsymbol{\varphi}_{\operatorname{S}}$ $\colon\mathbb{C}\times[0,\infty)\to\mathbb{C}^{n\times n}$ with $\boldsymbol{\varphi}_{\operatorname{S}}(s,\eta)$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}s^{2}{\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}+(1+\mathfrak{i}\eta){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}$,
$\mathcal{C}$ $\colon\mathbb{C}\to\mathbb{C}^{p\times n}$ with $\mathcal{C}(s)$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}}+s{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$.
For the minimization problem eq. 31, the gradients with respect to the optimization parameters are
$\frac{\partial}{\partial{\alpha}}\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)$ $=2\operatorname{Re}\bigg(s_{k}\operatorname{tr}\bigg({\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big(\boldsymbol{g}_{k}$ (33a)
@@
$\frac{\partial}{\partial{\beta}}\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)$ $=2\operatorname{Re}\bigg(s_{k}\operatorname{tr}\bigg({\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big(\boldsymbol{g}_{k}$ (33b)
@@
$\quad{}\times{}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big(\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big)^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{S}}(s_{k},\eta)^{-1}\bigg)\bigg).$
@@
It suffices to show eq. 33b since the proofs of eq. 33 follow analogously. Define $\mathcal{E}_{k}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{g}_{k}-\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\in\mathbb{C}^{p\times m}$ so that $\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)=\lVert\mathcal{E}_{k}\rVert_{\operatorname{F}}^{2}=\operatorname{tr}(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{E}_{k})$. Now, we consider an arbitrary perturbation $\Delta_{\beta}\in\mathbb{R}$, and expand the helper function $\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta+\Delta_{\beta})$ as
@@
$=\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}-s_{k}\Delta_{\beta}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}$
$\quad{}+{}\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2}).$
@@
$=\operatorname{tr}(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{E}_{k})+2\operatorname{Re}\Big(s_{k}\operatorname{tr}\Big(\mathcal{E}_{k}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\Big)\Big)\Delta_{\beta}$
$\quad{}+{}\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2})$
$=\mathcal{J}_{k,\operatorname{R}}(\alpha,\beta)+2\operatorname{Re}\Big(s_{k}\operatorname{tr}\Big({\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\mathcal{E}_{k}^{\mathsf{H}}\mathcal{C}(s_{k})\boldsymbol{\varphi}_{\operatorname{R}}(s_{k},\alpha,\beta)^{-1}\Big)\Big)\Delta_{\beta}$
$\quad{}+{}\mathcal{O}(\lVert\Delta_{\beta}\rVert_{\operatorname{F}}^{2}),$